The answer to each of the questions is a single digit integer ranging from 0 to 9.


 






     A very broad elevator is going up vertically with a constant acceleration of 2 m/s2. At the instant when its velocity is 4 m/s a ball is projected from the floor of the lift with a speed of 4 m/s relative to the floor at an elevation of 30o. The time taken by the ball to return the floor is  second. Find x. (g = 10 m/s2)



 

Grade:upto college level

4 Answers

Simran Bhatia
10 years ago

3

 

    Components of velocity of ball relative to lift are

        ux = 4 cos 30o = 2√3 m/s

    and    uy = 4 sin 30o = 2 m/s

        T = (2u_y)/12=u_y/6=2/6=1/3 s

         x = 3

Aditi Bardhan
5 years ago
Total acceleration=g+ 2m/s=12 m/sTherefore time of flight=(2usinΦ)/g=(2*4*1/2)/12 =1/3 s. Given [u=4m/s ; sin30=1/2]Hope this helps.... If you have any problem regarding this... Post it here.....
Gitanjali Rout
5 years ago
 A very broad elevator is going up vertically with a constant acceleration of 2 m/s2. At the instant when its velocity is 4 m/s a ball is projected from the floor of the lift with a speed of 4 m/s relative to the floor at an elevation of 30o. The time taken by the ball to return the floor is  second. Find x. (g = 10 m/s2)
Gitanjali Rout
5 years ago
otal acceleration=g+ 2m/s=12 m/sTherefore time of flight=(2usinΦ)/g=(2*4*1/2)/12 =1/3 s. Given [u=4m/s ; sin30=1/2]Hope this helps.... If you have any problem regarding this... Post it here.....nbk